OpenStudy (dancergirl45):

4. Ray needs help creating the second part of the coaster. Create a unique parabola in the pattern f(x) = ax2 + bx + c. Describe the direction of the parabola and determine the y-intercept and zeros.

OpenStudy (dancergirl45):

My parabola is f(x)= 3x^2+6x+15

OpenStudy (dancergirl45):

Here is the graph https://www.desmos.com/calculator/m9uxxepivl

OpenStudy (dancergirl45):

So would the answer be the parabola is going up because it is positive. And the y-int is 15

OpenStudy (dancergirl45):

@phi

OpenStudy (dancergirl45):

@Sachintha

OpenStudy (dancergirl45):

@AloneS

OpenStudy (dancergirl45):

I dont know about the zeros though

OpenStudy (phi):

ok on it being U shaped, with a y-intercept of 15 The zeros are where it crosses the x-axis. As you can see, it is well above the x-axis. You might think it has no zeros, but if we use complex numbers, we can find "zeros" You can find the complex roots (i.e. zeros) by using the quadratic formula or you can tweak your equation (instead of adding 15 at the end, use a negative number... that has the effect of lowering the whole shape downward)

OpenStudy (dancergirl45):

ok can we use -15

OpenStudy (phi):

yes, plot the new equation and see what you get.

OpenStudy (dancergirl45):

https://www.desmos.com/calculator/abydbjucdf

OpenStudy (dancergirl45):

so the new y-int in -15

OpenStudy (phi):

yes. You now have zeros, but they are not "nice" whole numbers. I would shift the curve up by 7 (I think). i.e. change the -15 to -8 and see what that looks like

OpenStudy (dancergirl45):

https://www.desmos.com/calculator/rnhnrvm29p

OpenStudy (dancergirl45):

so the new y-int is -8 and your zeros are -3 and 1

OpenStudy (phi):

ok, that is almost it. shift it down 1 (make the -8 a -9

OpenStudy (phi):

yes, I am trying to make the zeros -3 and 1. almost got it.

OpenStudy (dancergirl45):

https://www.desmos.com/calculator/psh5hg83x9

OpenStudy (dancergirl45):

So would the answer be the parabola is going up because it is positive. And the y-int is -9 while the zeros are -3 and 1.

OpenStudy (phi):

yes, if you "zoom in" you can see we go nice zeros, and a nice y-intercept (and a nice vertex, if you need that )

OpenStudy (phi):

the parabola is going up because the "a" coefficient on the x^2 term is positive.

OpenStudy (phi):

we say it is "convex" ( it's easier to remember "holds water" like a cup)

OpenStudy (phi):

if you know how to factor your equation you will find the zeros will be -3 and 1 you can use the equation to find its vertex is -b/(2a) or -6/(2*3) = -6/6= -1 i.e. the x value of the vertex is -1 (which we see from the graph, it is)

OpenStudy (dancergirl45):

ok thank you so much I have one more question could you please help me?

OpenStudy (dancergirl45):

It is basically about what you just said about the vertex

OpenStudy (phi):

?

OpenStudy (dancergirl45):

do you want me to open a new question

OpenStudy (phi):

no. what is the question ?

OpenStudy (dancergirl45):

5. The safety inspector notes that Ray also needs to plan for a vertical ladder through the center of the coaster's parabolic shape for access to the coaster to perform safety repairs. Find the vertex and the equation for the axis of symmetry of the parabola, showing your work, so Ray can include it in his coaster plan.

OpenStudy (phi):

Find the vertex and the equation for the axis of symmetry if you have an equation y = a x^2 + b x + c (and you do!) the x value of the vertex is at x= -b/(2a) can you do that calculation ? (I hope so, I posted it up above)

OpenStudy (dancergirl45):

yes you would get -6/6=-1

OpenStudy (phi):

ok, (but if you write this up, you should explain a is 3 and b is 6 ) the "axis of symmetry" is a line that divides the parabola in half. it is x=-1 so you have the axis of symmetry and the x value for the vertex. that last thing to find is the value of the vertex use your equation to find y when x is -1 can you do that? (btw, you know the answer by looking at your graph)

OpenStudy (dancergirl45):

so the axis of symmetry is the same as the x value of the vertex

OpenStudy (phi):

sort of. the answer is x= -1 but (unfortunately) that is a bit confusing. If we say that is the *equation of a line* x=-1 means a line where all the points on the line have an x value of -1 and the y value can be any number. if you did not know it was a line, then in algebra, x=-1 would mean the x variable has a value of -1, which is a different idea, so you should say: the axis of symmetry is the line x= -1 (then the reader will know that x= -1 is the equation of a line)

OpenStudy (phi):

but, can you find the y value of the vertex ? (using x=-1 in your equation for the parabola)

OpenStudy (dancergirl45):

https://www.desmos.com/calculator/os3npoyhof

OpenStudy (dancergirl45):

so the y value is 0

OpenStudy (phi):

yes, one of the zeros is +1 which means y=0 when x is 1 but you want to use x= -1

OpenStudy (dancergirl45):

ok so it is -12

OpenStudy (phi):

and you might want to practice doing it by hand, because they will expect you to know how. yes, the vertex is at (-1, -12) can you find that point on your graph of the parabola?

OpenStudy (dancergirl45):

https://www.desmos.com/calculator/4dy7eewmec

OpenStudy (dancergirl45):

yes i found the point

OpenStudy (dancergirl45):

Thank you so much not only did you help me get the answer, now I know how to do these types of problems :)